# Janus-Q: End-to-End Event-Driven Trading via Hierarchical-Gated Reward Modeling

### v1

@@ 1. Introduction
Financial markets have long been studied through the lens of time-series modeling, where asset prices or returns are predicted directly from historical numerical signals such as prices, volumes, and technical indicators (Zhou et al., 2025a; Zhang et al., 2025b). Despite decades of methodological advances, purely numerical forecasting remains fundamentally challenging due to severe noise, non-stationarity, and frequent regime shifts in real-world markets (Fama, 1970; Lo, 2004). These limitations have motivated growing interest in incorporating alternative information sources, including financial news, corporate disclosures, and social media, into learning-based trading systems (Zhou et al., 2025b; Wang et al., 2026; Zhou et al., 2024).
@@
Although most market-moving events are communicated through unstructured financial news, existing learning-based trading systems have yet to fully exploit event-level information. The dominant paradigm embeds textual news and fuses it with historical price sequences under numerical forecasting objectives (Kong et al., 2025a; Zhang et al., 2025c). Under this formulation, textual information is typically treated as an auxiliary modality, rather than as a primary driver of trading decisions (Li et al., 2024a; Dong et al., 2024; Saqur et al., 2024). As a result, learned policies are often dominated by recent price dynamics, even when those dynamics contradict the semantic implications of newly released events. Figure 1 illustrates this structural mismatch between time-series–driven and event-driven trading paradigms. This limitation stems primarily from the lack of structured supervision at the event level. While prior work has explored sentiment labels or coarse market reactions (Tetlock, 2007b; Wang et al., 2025b; Li et al., 2025), two fundamental challenges remain largely unaddressed.
@@
Challenge 2 (Ch2): Misalignment Between Semantic Reasoning and Market Reality. Beyond data limitations, aligning language model reasoning with empirical market behavior remains nontrivial. Although large language models (LLM) can generate fluent and plausible interpretations of financial news, their semantic judgments are not inherently grounded in realized market outcomes (Zhang et al., 2024; Zhou et al., 2025c). In practice, semantic polarity does not map linearly to price movements: apparently positive announcements may trigger corrections due to priced-in expectations, while negative news can be absorbed by the market and followed by stabilization or rebound. Consequently, purely supervised learning risks capturing superficial correlations, whereas profit-driven optimization alone often induces spurious strategies that exploit short-term noise (Liu et al., 2024).
@@
CSMAR (Wang et al., 2025a) General 13 / CN
@@ 2.1. Event-Driven Modeling
Early studies on event-driven market analysis originate from econometric and statistical finance, where discrete corporate or macroeconomic events are treated as exogenous shocks and analyzed through abnormal return dynamics around predefined event windows. The classical event study methodology formalized this process by estimating expected returns from factor models and attributing deviations to the information content of events, establishing a standard toolkit for measuring market reactions (MacKinlay, 1997b). As textual disclosures and news became increasingly accessible, subsequent research incorporated qualitative information into event analysis, showing that media tone and sentiment convey economically meaningful signals that explain and predict short-horizon market movements beyond purely price-based features (Tetlock, 2007a). Building on this line, machine learning (ML) and deep learning (DL) approaches began to integrate textual information with market data, shifting from handcrafted sentiment indicators toward learning event semantics directly from text. Ding et al. (Ding et al., 2015) extract structured events from news and learn dense event representations to model both short-term and long-term effects on stock movements, with subsequent work further enhancing event representations by incorporating richer contextual and temporal signals (Ding et al., 2016). More recently, Wang et al. (Wang et al., 2025a) introduce StockMem, an event-reflection memory framework that organizes news into structured events and leverages their temporal evolution to retrieve analogous historical scenarios, supporting more explainable stock movement prediction. However, despite richer event representations, existing pipelines remain prediction‑oriented and treat events primarily as inputs, without explicitly modeling event‑level properties such as historical context or relative importance to support decision‑oriented trading objectives. Moreover, the field still lacks a comprehensive, standardized dataset for trading. Following this direction, we propose an event‑centric dataset that integrates event categorization, sentiment annotation, and CAR‑based evaluation of event impacts.
@@ 4.4. Baseline
$\bullet$ Financial Domain-Specific LLM. We evaluate several open-source language models pre-trained or fine-tuned on financial corpora, including FinMA (Xie et al., 2023), DISC-FinLLM (Chen et al., 2023), and Stock-Chain (Li et al., 2024b). These models are designed to capture financial semantics and domain-specific patterns, enabling direct application to market analysis and trading-related evaluation.
$\bullet$ General-Purpose LLM. We further include widely used general-purpose language models, including QwQ-32B (Yang et al., 2025a), Claude-3-Haiku (Rahman et al., 2024), GPT-4o-mini (Hurst et al., 2024), DeepSeek-v3.1-nex-n1 (Cai et al., 2025), Grok-3-mini-beta (Jiang et al., 2025), Qwen2.5-7B (Team and others, 2024), and Gemini-2.5-flash (Comanici et al., 2025), which are not specifically tuned for financial tasks.

### v2

# arXiv:2602.19919v2[cs.CL] 27 Feb 2026

@@ 1. Introduction
Financial markets have long been studied through the lens of time-series modeling, where asset prices or returns are predicted directly from historical numerical signals such as prices, volumes, and technical indicators (Zhou et al., 2025b; Zhang et al., 2025b). Despite decades of methodological advances, purely numerical forecasting remains fundamentally challenging due to severe noise, non-stationarity, and frequent regime shifts in real-world markets (Fama, 1970; Lo, 2004). These limitations have motivated growing interest in incorporating alternative information sources, including financial news, corporate disclosures, and social media, into learning-based trading systems (Zhou et al., 2025c; Wang et al., 2026; Zhou et al., 2024).
@@
Although most market-moving events are communicated through unstructured financial news, existing learning-based trading systems have yet to fully exploit event-level information. The dominant paradigm embeds textual news and fuses it with historical price sequences under numerical forecasting objectives (Kong et al., 2025a; Zhang et al., 2025c). Under this formulation, textual information is typically treated as an auxiliary modality, rather than as a primary driver of trading decisions (Li et al., 2024b; Dong et al., 2024; Saqur et al., 2024). As a result, learned policies are often dominated by recent price dynamics, even when those dynamics contradict the semantic implications of newly released events. Figure 1 illustrates this structural mismatch between time-series–driven and event-driven trading paradigms. This limitation stems primarily from the lack of structured supervision at the event level. While prior work has explored sentiment labels or coarse market reactions (Tetlock, 2007a; Wang et al., 2025a; Li et al., 2025), two fundamental challenges remain largely unaddressed.
@@
Challenge 2 (Ch2): Misalignment Between Semantic Reasoning and Market Reality. Beyond data limitations, aligning language model reasoning with empirical market behavior remains nontrivial. Although large language models (LLM) can generate fluent and plausible interpretations of financial news, their semantic judgments are not inherently grounded in realized market outcomes (Zhang et al., 2024; Zhou et al., 2025a). In practice, semantic polarity does not map linearly to price movements: apparently positive announcements may trigger corrections due to priced-in expectations, while negative news can be absorbed by the market and followed by stabilization or rebound. Consequently, purely supervised learning risks capturing superficial correlations, whereas profit-driven optimization alone often induces spurious strategies that exploit short-term noise (Liu et al., 2024).
@@
CSMAR (Wang et al., 2025b) General 13 / CN
@@ 2.1. Event-Driven Modeling
Early studies on event-driven market analysis originate from econometric and statistical finance, where discrete corporate or macroeconomic events are treated as exogenous shocks and analyzed through abnormal return dynamics around predefined event windows. The classical event study methodology formalized this process by estimating expected returns from factor models and attributing deviations to the information content of events, establishing a standard toolkit for measuring market reactions (MacKinlay, 1997b). As textual disclosures and news became increasingly accessible, subsequent research incorporated qualitative information into event analysis, showing that media tone and sentiment convey economically meaningful signals that explain and predict short-horizon market movements beyond purely price-based features (Tetlock, 2007b). Building on this line, machine learning (ML) and deep learning (DL) approaches began to integrate textual information with market data, shifting from handcrafted sentiment indicators toward learning event semantics directly from text. Ding et al. (Ding et al., 2015) extract structured events from news and learn dense event representations to model both short-term and long-term effects on stock movements, with subsequent work further enhancing event representations by incorporating richer contextual and temporal signals (Ding et al., 2016). More recently, Wang et al. (Wang et al., 2025b) introduce StockMem, an event-reflection memory framework that organizes news into structured events and leverages their temporal evolution to retrieve analogous historical scenarios, supporting more explainable stock movement prediction. However, despite richer event representations, existing pipelines remain prediction‑oriented and treat events primarily as inputs, without explicitly modeling event‑level properties such as historical context or relative importance to support decision‑oriented trading objectives. Moreover, the field still lacks a comprehensive, standardized dataset for trading. Following this direction, we propose an event‑centric dataset that integrates event categorization, sentiment annotation, and CAR‑based evaluation of event impacts.
@@ 4.4. Baseline
$\bullet$ Financial Domain-Specific LLM. We evaluate several open-source language models pre-trained or fine-tuned on financial corpora, including FinMA (Xie et al., 2023), DISC-FinLLM (Chen et al., 2023), and Stock-Chain (Li et al., 2024a). These models are designed to capture financial semantics and domain-specific patterns, enabling direct application to market analysis and trading-related evaluation.
$\bullet$ General-Purpose LLM. We further include widely used general-purpose language models, including QwQ-32B (Yang et al., 2025a), Claude-3-Haiku (Rahman et al., 2024), GPT-4o-mini (Hurst et al., 2024), DeepSeek-v3.1-nex-n1 (Cai et al., 2025), Grok-3-mini-beta (Jiang et al., 2025), Qwen2.5-7B (Team et al., 2024), and Gemini-2.5-flash (Comanici et al., 2025), which are not specifically tuned for financial tasks.